OpenStudy (anonymous):

verify the identity cos(x+ pi/2) = - sin x

jimthompson5910 (jim_thompson5910):

hint: cos(A+B) = cos(A)cos(B) - sin(A)sin(B)

OpenStudy (dan815):

^---- hint too big

jimthompson5910 (jim_thompson5910):

what do you mean dan815?

OpenStudy (jhannybean):

\[\large cos (\color{red}A +\color{blue}B) = cos(\color{red}A)cos(\color{blue}B) - sin(\color{red}A)sin(\color{blue}B)\]\[\large cos(\color{red}x+ \color{blue}{\pi/2})= \]

OpenStudy (jhannybean):

I love these formulas cause you can apply COLORS!!!

OpenStudy (dan815):